# Resilience for Regular Path Queries: Towards a Complexity Classification

### v2

@@ 5.4. Reduced Languages of Star-Free Languages are Star-Free
We now start proving the statement. Let us first show that (call this $\dagger$) if both $\alpha^{\prime}$ and $\delta^{\prime}$ are empty then the result holds. Indeed, in that case we have $x\beta^{\prime}\in L^{\prime}$ and $\gamma^{\prime}x\in L^{\prime}$. By ($\star$) (first two points) we have that $\beta^{\prime}$ and $\gamma^{\prime}$ are both nonempty. By Section 5.6 we deduce from $x\beta^{\prime}\in L^{\prime}$ that $xe\beta^{\prime}\in L^{\prime}$, and we deduce from $\gamma^{\prime}x\in L^{\prime}$ that $\gamma^{\prime}ex\in L^{\prime}$. Now, take $e$ as a candidate choice of body, and $\alpha=x$, $\beta=\beta^{\prime}$, $\gamma=\gamma^{\prime}$ and $\delta=x$, as a candidate choice of legs: all of these are non-empty. We have $\alpha e\beta\in L^{\prime}$ and $\gamma e\delta\in L^{\prime}$. So, considering the word $\alpha e\delta=xex$, there are two possibilities: either $xex\notin L^{\prime}$ or $xex\in L^{\prime}$. In the first case, we conclude with our choice of body and legs that $L^{\prime}$ is four-legged, which allows us to conclude. In the second case, we have $xx\in L$ because $e$ is neutral for $L$. Further, we have $xx\in L^{\prime}$ because the strict infixes of $xx$ are the empty word and $x$ which are both strict infixes of the word $x\beta^{\prime}\in L^{\prime}$: we conclude because $L^{\prime}$ is reduced. The fact that $xx\in L^{\prime}$ allows us to conclude: indeed, we can then do the exact same reduction from vertex cover as in Proposition 4.1 (noticing that we do not use other letters than ’$x$’).
The same reasoning shows that (call this $\dagger^{\prime}$) if both $\beta^{\prime}$ and $\gamma^{\prime}$ are empty then the result holds. Indeed, in this case we have $\alpha^{\prime}x\in L^{\prime}$ and $x\delta^{\prime}\in L^{\prime}$, and we have $\alpha^{\prime}$ and $\delta^{\prime}$ both nonempty by ($\star$) (first two points). Applying Section 5.6 to each gives us that $\alpha^{\prime}ex\in L^{\prime}$ and $xe\delta^{\prime}\in L^{\prime}$, and again we conclude that $L^{\prime}$ is four-legged or that $xex\in L^{\prime}$ from which we show $xx\in L^{\prime}$ in the same way.
We have ruled out the case of $\alpha^{\prime}$ and $\delta^{\prime}$ being both empty, and of $\beta^{\prime}$ and $\gamma^{\prime}$ being both empty. Combining with ()($\star$)( ⋆ ), one gets that either $\alpha^{\prime}$ and $\gamma^{\prime}$ are both non-empty, or $\beta^{\prime}$ and $\delta^{\prime}$ are both non-empty. Indeed, assume that the first term in the disjunction is false, i.e., that it is not the case that $\alpha^{\prime}$ and $\gamma^{\prime}$ are both non-empty. If we have $\alpha^{\prime}=\epsilon$, then $\delta^{\prime}$ is non-empty from ($\dagger$), and $\beta^{\prime}$ is non-empty by ()($\star$)( ⋆ ) (first point). If we have $\gamma^{\prime}=\epsilon$, then $\beta^{\prime}$ is non-empty by ($\dagger^{\prime}$), and $\delta^{\prime}$ is non-empty by ()($\star$)( ⋆ ) (second point).
@@
  • Assume that $\alpha^{\prime}\neq\epsilon$ and $\gamma^{\prime}\neq\epsilon$. We have $\alpha^{\prime}x\beta^{\prime}\in L^{\prime}$ so by Section 5.6 we have $\alpha^{\prime}ex\beta^{\prime}\in L^{\prime}$, and we have $\gamma^{\prime}x\delta^{\prime}\in L^{\prime}$ so by Section 5.6 we have $\gamma^{\prime}ex\delta^{\prime}\in L^{\prime}$. We then claim that $\alpha^{\prime}ex\delta^{\prime}\notin L^{\prime}$, implying that $L^{\prime}$ is four-legged (taking $e$ as the body, $\alpha=\alpha^{\prime}$, $\beta=x\beta^{\prime}$, $\gamma=\gamma^{\prime}$ and $\delta=x\delta^{\prime}$, which are all not empty). Indeed, assume by way of contradiction that $\alpha^{\prime}ex\delta^{\prime}\in L^{\prime}$. We would then have $\alpha^{\prime}ex\delta^{\prime}\in L$, so $\alpha^{\prime}x\delta^{\prime}\in L$ (as $e$ is neutral). As $\alpha^{\prime}x\delta^{\prime}\notin L^{\prime}$ by assumption, there exists a strict infix $\tau$ of $\alpha^{\prime}x\delta^{\prime}$ that is in $L^{\prime}$. Now, adding $e$ to $\tau$ would give a strict infix $\tau^{\prime}$ of $\alpha^{\prime}ex\delta^{\prime}$, with $\tau^{\prime}\in L$ because $\tau\in L$ and $e$ is neutral for $L$, a contradiction because $L^{\prime}$ is reduced.
  • Assume that $\beta^{\prime}\neq\epsilon$ and $\delta^{\prime}\neq\epsilon$. The reasoning is similar: using Section 5.6 we have $\alpha^{\prime}xe\beta^{\prime}\in L^{\prime}$ and $\gamma^{\prime}xe\delta^{\prime}\in L^{\prime}$ and we obtain that $L^{\prime}$ is four-legged by taking $e$ as the body, $\alpha=\alpha^{\prime}x$, $\beta=\beta^{\prime}$, $\gamma=\gamma^{\prime}x$ and $\delta=\delta^{\prime}$.
@@
For completeness we give here examples of languages for which only one of the two cases in the statement of Section 5.6 holds. Let $L_{1}=e^{*}be^{*}ce^{*}|e^{*}de^{*}fe^{*}$ (for which $e$ is neutral), with $\mathtt{reduce}(L_{1})=be^{*}c|de^{*}f$. Then $\mathtt{reduce}(L_{1})$ is four-legged (taking $x=e$, $\alpha=b$, $\beta=c$, $\gamma=d$, $\delta=f$), hence $\mathtt{reduce}(L_{1})$ is not local, but it does not have a word of the form $xx$ for $x\in\Sigma$. Now take $L_{2}=e^{*}(a|c)e^{*}(a|d)e^{*}$, with $\mathtt{reduce}(L_{2})=(a|c)e^{*}(a|d)$. Then $\mathtt{reduce}(L_{2})$ is not local (taking $x=a$, $\alpha=c$, $\beta=\epsilon$, $\gamma=\epsilon$, $\delta=d$), it contains the word $aa$, but one can check that it is not four-legged.
We can then use Section 5.6 to prove Section 5.6:
Proof of Section 5.6.
If $\mathtt{reduce}(L)$ is local then $\mathtt{RES}_{\mathrm{bag}}(L)$ is PTIME by Theorem 3.3. Otherwise, by Section 5.6 either $\mathtt{reduce}(L)$ is four-legged, in which case $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard by Section 5.2, or $aa\in\mathtt{reduce}(L)$ for $a\in\Sigma$. In this case one can easily adapt the proof of Proposition 4.1: one can check that the reduction only builds instances using the letter $a$, and the restriction of $\mathtt{reduce}(L)$ to words using only the letter $a$ must be exactly $aa$ given that $\mathtt{reduce}(L)$ is reduced. This concludes the proof. ∎
The dichotomy above relies on the presence of a neutral letter, so it does not apply to all languages. In particular, it does not apply at all to finite languages, so we focus on them from now on.
@@ 6. Hardness of Finite Languages Having Repeated Letters
We now conclude the presentation of our main hardness results by showing that resilience is hard for a class of languages that generalizes $aa$ (Proposition 4.1). We say that a word $\alpha$ has a repeated letter if we have $\alpha=\beta a\gamma a\delta$ for some $a\in\Sigma$ and $\beta,\gamma,\delta\in\Sigma^{*}$. We show that, for reduced and finite languages, resilience is NP-hard whenever the language contains a word with a repeated letter:
{toappendix}
6.1. Proof of Section 6.1: Hardness with Repeated Letters
{theoremrep}
[Hardness with repeated letters] Let $L$ be a reduced finite language that contains a word with a repeated letter. Then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-complete.
{proofsketch}
The proof is rather technical. We use the finiteness of $L$ to select a word with a repeated letter of the form $\beta a\gamma a\delta$ where the gap between repeated letters is maximal. A long case distinction then allows us to deduce that either $L$ is four-legged (in which case the result follows from Section 5.2), or we can show that one of several gadgets applies.
{toappendix}
The proof of Section 6.1 is technical, so we sketch it here before giving the details in the rest of the section. We first make a preliminary observation: the tractability of resilience is preserved under the mirror operation on regular languages. We then start from a word featuring a repeated letter, and pick it to optimize certain criteria (maximizing the gap between the repeated letters, and then maximizing the total length of the word): we call this a maximal-gap word, and the existence of such words is the only place where we use the fact that $L$ is finite.
@@
We first make a straightforward observation on the mirror of languages. The mirror of a word $\alpha=a_{1}\cdots a_{n}\in\Sigma^{*}$ is $\alpha^{\mathrm{R}}=a_{n}\cdots a_{1}\in\Sigma^{*}$ (in particular $\epsilon^{\mathrm{R}}=\epsilon$). The mirror of a language $L\subseteq\Sigma^{*}$ is $L^{\mathrm{R}}\coloneq\{\alpha^{\mathrm{R}}\mid\alpha\in L\}$. Observe that $(L^{\mathrm{R}})^{\mathrm{R}}=L$. It is immediate by symmetry that a language $L$ is reduced iff $L^{\mathrm{R}}$ is reduced. Further, it is also immediate that a language and its mirror have the same complexity for resilience:
{propositionrep}
@@
The point of this claim will be to eliminate symmetric cases, up to replacing the language by its mirror. We only use the claim about set semantics in the sequel (given that we will show hardness of resilience in set semantics, from which hardness for bag semantics follows): the claim for bag semantics in Section 6.1 is given just for completeness.
Proof of Section 6.1.
@@
Let $L$ be a reduced finite language that contains a word with a repeated letter, and let us show that $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard, to establish Section 6.1. We will pick a word with a repeated letter for the rest of the proof, but we will choose such a word carefully. Here is the precise notion we will need:
@@
It is clear that any finite language that contains a word with a repeated letter must have maximal-gap words: this is the only place where we use the finiteness of the language. Hence, in the rest of the proof, we fix $\alpha=\beta a\gamma a\delta$ to be an arbitrary such word. We will do case disjunctions on this word in the rest of the proof.
@@
{claimrep}
Let $L^{\prime}$ be a reduced language containing a maximal-gap word $\beta a\gamma a\delta$ such that $\beta\neq\epsilon$ and $\delta\neq\epsilon$. Then $L^{\prime}$ is four-legged.
@@
Consider the words $\alpha=(\beta a\gamma)a\delta$ and $\alpha=\beta a(\gamma a\delta)$. Note that $\beta a\gamma$, $\delta$, $\beta$, and $\gamma a\delta$ are all nonempty, so they can be used as legs. So consider the word $\alpha^{\prime}=(\beta a\gamma)a(\gamma a\delta)$. The word $\alpha^{\prime}$ contains two $a$’s separated by $\gamma a\gamma$, which is longer than $\gamma$. Hence, $\alpha^{\prime}$ is not in in $L^{\prime}$, as it would contradict the fact that $\alpha$ is maximal-gap (the first criterion of the definition of maximal-gap suffices). So we have a witness of the fact that $L^{\prime}$ is four-legged. ∎
If $\beta$ and $\delta$ are both nonempty, then applying Section 6.1 and Section 5.2 with $L^{\prime}\coloneq L$ concludes. Hence, in the rest of the argument, we assume that one of $\beta$ and $\delta$ is empty. Further, without loss of generality, up to replacing $L$ by $L^{\mathrm{R}}$ using Section 6.1, we assume that $\beta=\epsilon$. Hence, in the rest of the proof, we can assume that $\beta=\epsilon$: we have a maximal-gap word $\alpha=a\gamma a\delta\in L$ and this word is maximal-gap.
@@
{lemmarep}
Let $L^{\prime}$ be a reduced language containing a word $a\gamma a\delta$ where no infix of $\gamma a\gamma$ is in $L^{\prime}$. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is NP-hard.
Figure 8. Gadget for the proof of Section 6.1 in the case when $\delta=\epsilon$ (left), and condensed hypergraph of matches (right).
Figure 9. Gadget for Section 6.1 in the case where $\delta\neq\epsilon$ (left), and condensed hypergraph of matches (right).
@@
In the case where $\delta=\epsilon$, we use the completed gadget pictured on Figure 8 (left), where the edge labeled $\gamma$ stands for a (possibly empty) path. Note that, no matter what $\gamma$ stands for, there will be no facts having $t_{u}$ and $t_{v}$ as their head (except the two facts of the gadget completion, pictured in orange), so indeed it is a pre-gadget. The condensed hypergraph of matches is shown on Figure 8 (right), and it is an odd chain, so indeed the instance pictured on Figure 8 (left) is (the completion of) a gadget and we can conclude by Section 4.8. Note how we use the assumption that the matches on the completed gadget are the ones indicated: directed paths including two $a$’s must include $a\gamma a$ as an infix and so they are dominated, and directed paths including only one $a$ cannot be a match of the query.
In the case where $\delta\neq\epsilon$, we use the (completed) gadget pictured on Figure 9 (left), again with the edge labeled $\gamma$ standing for a possibly empty path. Again, no matter what $\gamma$ stands for, there will be no facts having $t_{u}$ and $t_{v}$ as head (except for the facts added in the completion), so again we have a pre-gadget. The condensed hypergraph of matches is shown on Figure 9 (right). Now, to show that it is actually like this, first observe that the facts of the $\delta$-paths are always dominated by the preceding $a$: there is no match of the query within $\delta$ (as it is a strict infix of $a\gamma a\delta$), so all matches using $\delta$ must use the preceding $a$, and we can remove the facts of $\delta$-paths in the condensation via the node domination rule. We are then essentially back to the preceding case. More precisely, again, no directed path containing only one copy of the $a$-edges of the gadget can be a word of the language, because these paths are either infixes of $\gamma a\delta$ (hence strict infixes of $a\gamma a\delta\in L$) or they are infixes of $\gamma a\gamma$ and so they are not in the language by assumption. Further, for any two consecutive $a$’s in the gadget, there is a $a\gamma a\delta$ match containing them. Further, after observing that $\delta$ facts are eliminated by node domination, we note that all query matches that are infixes of a word of the form $a(\gamma a)^{*}$ must contain two consecutive $a$’s, so they are subsumed by the matches corresponding to $a\gamma a\delta$, which after node domination contain just two consecutive $a$’s and the $\gamma$ between them. (The facts corresponding to $\gamma$ are then eliminated by node domination.)
Hence, we indeed have a condensed hypergraph of matches corresponding to Figure 9 (right), and we can conclude by Section 4.8. ∎
@@
{claim}
Let $L^{\prime}$ be a reduced language containing the word $\alpha=a\gamma a\delta$ and containing an infix of $\gamma a\gamma$. Then $L^{\prime}$ contains a word of the form $\alpha^{\prime}=\gamma_{1}a\gamma_{2}$ where $\gamma_{1}$ is a non-empty suffix of $\gamma$ and $\gamma_{2}$ is a non-empty prefix of $\gamma$.
@@
Consider the infix of $\gamma a\gamma$ which exists by hypothesis. This infix must include the middle $a$, as otherwise it is an infix of $\gamma$, which is a strict infix of $\alpha\in L^{\prime}$, and this contradicts the assumption that $L^{\prime}$ is reduced.

### v3

@@ 5.2. Implications of Hardness for Four-Legged Languages
We now start proving the statement. Let us first show that (call this $\dagger$) if both $\alpha^{\prime}$ and $\delta^{\prime}$ are empty then the result holds. Indeed, in that case we have $x\beta^{\prime}\in L^{\prime}$ and $\gamma^{\prime}x\in L^{\prime}$. By ($\star$) (first two points) we have that $\beta^{\prime}$ and $\gamma^{\prime}$ are both non-empty. By 5.9 we deduce from $x\beta^{\prime}\in L^{\prime}$ that $xe\beta^{\prime}\in L^{\prime}$, and we deduce from $\gamma^{\prime}x\in L^{\prime}$ that $\gamma^{\prime}ex\in L^{\prime}$. Now, take $e$ as a candidate choice of body, and $\alpha=x$, $\beta=\beta^{\prime}$, $\gamma=\gamma^{\prime}$ and $\delta=x$, as a candidate choice of legs: all of these are non-empty. We have $\alpha e\beta\in L^{\prime}$ and $\gamma e\delta\in L^{\prime}$. So, considering the word $\alpha e\delta=xex$, there are two possibilities: either $xex\notin L^{\prime}$ or $xex\in L^{\prime}$. In the first case, we conclude with our choice of body and legs that $L^{\prime}$ is four-legged, which allows us to conclude. In the second case, we have $xx\in L$ because $e$ is neutral for $L$. Further, we have $xx\in L^{\prime}$ because the strict infixes of $xx$ are the empty word and $x$ which are both strict infixes of the word $x\beta^{\prime}\in L^{\prime}$: we conclude because $L^{\prime}$ is infix-free. The fact that $xx\in L^{\prime}$ allows us to conclude: indeed, we can then do the exact same reduction from vertex cover as in Proposition 4.1 (noticing that we do not use other letters than ’$x$’).
The same reasoning shows that (call this $\dagger^{\prime}$) if both $\beta^{\prime}$ and $\gamma^{\prime}$ are empty then the result holds. Indeed, in this case we have $\alpha^{\prime}x\in L^{\prime}$ and $x\delta^{\prime}\in L^{\prime}$, and we have $\alpha^{\prime}$ and $\delta^{\prime}$ both non-empty by ($\star$) (first two points). Applying 5.9 to each gives us that $\alpha^{\prime}ex\in L^{\prime}$ and $xe\delta^{\prime}\in L^{\prime}$, and again we conclude that $L^{\prime}$ is four-legged or that $xex\in L^{\prime}$ from which we show $xx\in L^{\prime}$ in the same way.
We have ruled out the case of $\alpha^{\prime}$ and $\delta^{\prime}$ being both empty, and of $\beta^{\prime}$ and $\gamma^{\prime}$ being both empty. Combining with $(\star)$, one gets that either $\alpha^{\prime}$ and $\gamma^{\prime}$ are both non-empty, or $\beta^{\prime}$ and $\delta^{\prime}$ are both non-empty. Indeed, assume that the first term in the disjunction is false, i.e., that it is not the case that $\alpha^{\prime}$ and $\gamma^{\prime}$ are both non-empty. If we have $\alpha^{\prime}=\epsilon$, then $\delta^{\prime}$ is non-empty from ($\dagger$), and $\beta^{\prime}$ is non-empty by $(\star)$ (first point). If we have $\gamma^{\prime}=\epsilon$, then $\beta^{\prime}$ is non-empty by ($\dagger^{\prime}$), and $\delta^{\prime}$ is non-empty by $(\star)$ (second point).
@@
  • Assume that $\alpha^{\prime}\neq\epsilon$ and $\gamma^{\prime}\neq\epsilon$. We have $\alpha^{\prime}x\beta^{\prime}\in L^{\prime}$ so by 5.9 we have $\alpha^{\prime}ex\beta^{\prime}\in L^{\prime}$, and we have $\gamma^{\prime}x\delta^{\prime}\in L^{\prime}$ so by 5.9 we have $\gamma^{\prime}ex\delta^{\prime}\in L^{\prime}$. We then claim that $\alpha^{\prime}ex\delta^{\prime}\notin L^{\prime}$, implying that $L^{\prime}$ is four-legged (taking $e$ as the body, $\alpha=\alpha^{\prime}$, $\beta=x\beta^{\prime}$, $\gamma=\gamma^{\prime}$ and $\delta=x\delta^{\prime}$, which are all not empty). Indeed, assume by way of contradiction that $\alpha^{\prime}ex\delta^{\prime}\in L^{\prime}$. We would then have $\alpha^{\prime}ex\delta^{\prime}\in L$, so $\alpha^{\prime}x\delta^{\prime}\in L$ (as $e$ is neutral). As $\alpha^{\prime}x\delta^{\prime}\notin L^{\prime}$ by assumption, there exists a strict infix $\tau$ of $\alpha^{\prime}x\delta^{\prime}$ that is in $L^{\prime}$. Now, adding $e$ to $\tau$ would give a strict infix $\tau^{\prime}$ of $\alpha^{\prime}ex\delta^{\prime}$, with $\tau^{\prime}\in L$ because $\tau\in L$ and $e$ is neutral for $L$, a contradiction because $L^{\prime}$ is infix-free.
  • Assume that $\beta^{\prime}\neq\epsilon$ and $\delta^{\prime}\neq\epsilon$. The reasoning is similar: using 5.9 we have $\alpha^{\prime}xe\beta^{\prime}\in L^{\prime}$ and $\gamma^{\prime}xe\delta^{\prime}\in L^{\prime}$ and we obtain that $L^{\prime}$ is four-legged by taking $e$ as the body, $\alpha=\alpha^{\prime}x$, $\beta=\beta^{\prime}$, $\gamma=\gamma^{\prime}x$ and $\delta=\delta^{\prime}$.
@@
For completeness we give here examples of languages for which only one of the two cases in the statement of Lemma 5.8 holds. Let $L_{1}=e^{*}be^{*}ce^{*}|e^{*}de^{*}fe^{*}$ (for which $e$ is neutral), with $\mathtt{IF}(L_{1})=be^{*}c|de^{*}f$. Then $\mathtt{IF}(L_{1})$ is four-legged (taking $x=e$, $\alpha=b$, $\beta=c$, $\gamma=d$, $\delta=f$), hence $\mathtt{IF}(L_{1})$ is not local, but it does not have a word of the form $xx$ for $x\in\Sigma$. Now take $L_{2}=e^{*}(a|c)e^{*}(a|d)e^{*}$, with $\mathtt{IF}(L_{2})=(a|c)e^{*}(a|d)$. Then $\mathtt{IF}(L_{2})$ is not local (taking $x=a$, $\alpha=c$, $\beta=\epsilon$, $\gamma=\epsilon$, $\delta=d$), it contains the word $aa$, but one can check that it is not four-legged.
We can then use Lemma 5.8 to prove our complexity characterization of resilience for languages featuring a neutral letter (Proposition 5.7):
Proof of Proposition 5.7.
If $\mathtt{IF}(L)$ is local then $\mathtt{RES}_{\mathrm{bag}}(L)$ is PTIME by Theorem 3.13. Otherwise, by Lemma 5.8 either $\mathtt{IF}(L)$ is four-legged, in which case $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard by Theorem 5.3, or $aa\in\mathtt{IF}(L)$ for $a\in\Sigma$. In this case one can easily adapt the proof of Proposition 4.1: one can check that the reduction only builds instances using the letter $a$, and the restriction of $\mathtt{IF}(L)$ to words using only the letter $a$ must be exactly $aa$ given that $\mathtt{IF}(L)$ is infix-free. This concludes the proof. ∎
The dichotomy above relies on the presence of a neutral letter, so it does not apply to all languages. In particular, it does not apply at all to finite languages. This motivates further study of finite languages, and in fact all the languages studied in the remainder of the paper will be finite languages.
@@ 6. Hardness of Finite Languages Having Repeated Letters
We now conclude the presentation of our main hardness results by showing that resilience is hard for a class of languages that generalizes $aa$ (Proposition 4.1). We say that a word $\alpha$ has a repeated letter if we have $\alpha=\beta a\gamma a\delta$ for some $a\in\Sigma$ and $\beta,\gamma,\delta\in\Sigma^{*}$. We show that, for infix-free finite languages, resilience is NP-hard whenever the language contains a word with a repeated letter:
Theorem 6.1 (Hardness with repeated letters).
Let $L$ be an infix-free finite language that contains a word with a repeated letter. Then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-complete.
Before proving Theorem 6.1, we discuss its relation with existing work and with the other results we have obtained so far.
Connections to UCQs and self-joins.
Observe that finite languages $L$ correspond to the case of RPQs that are in fact unions of conjunctive queries (UCQs). The UCQs obtained from RPQs have a specific form: they are posed on an arity-two signature corresponding to the alphabet $\Sigma$, and their constituent conjunctive queries (CQs) are directed paths corresponding to words of the language $L$. In this sense, repeated letters in words correspond to self-joins in CQs. Thus, Theorem 6.1 implies that, for those UCQs that correspond to finite RPQs, the resilience problem is NP-hard unless every CQ is self-join-free. This contrasts with the setting of general UCQs, where resilience is sometimes tractable even in the presence of self-joins (DBLP:conf/pods/FreireGIM20, ). Specifically, in that setting, the following Boolean CQ on an arity-two signature is known to be in PTIME (DBLP:conf/pods/FreireGIM20, ) despite having two self-joins: Q:A(x),B(x),R(x,y),R(z,y),A(z),C(z)Q{\,:\!\!-\,}A(x),B(x),R(x,y),R(z,y),A(z),C(z).
Connections to other language classes.
We now explain how Theorem 6.1 relates to the languages discussed so far. We first perform a sanity check: such languages are never local.
Lemma 6.2.
Proof.
As $L$ is finite and contains a word with a repeated letter, we can take a word with a repeated letter whose length is maximum among words of $L$ with a repeated letter. Let $\alpha=\beta a\gamma a\delta\in L$ be such a word. Assume by contradiction that $L$ is local. Then, by Proposition 3.5, $L$ is letter-Cartesian. Applying the definition to $\alpha$ and $\alpha$, we must have that the following word is also in $L$:
$\beta a\gamma a\gamma a\delta\in L.$
Note that the finiteness hypothesis is essential for Lemma 6.2, e.g., $L=ax^{*}b$ contains words with repeated letters but it is local (see Example 3.2). Further notice that there is no converse to Lemma 6.2: there are finite infix-free non-local languages that feature no words with repeated letters and for which resilience is hard. This is for instance the case of the four-legged language $axb|cxd$ of Proposition 4.13. Thus, the languages covered by Theorem 6.1 and Theorem 5.3 are incomparable: they each generalize Proposition 4.1 and Proposition 4.13, respectively, without implying the other result. (However, the proof of Theorem 6.1 relies on Theorem 5.3.) These two results also do not cover all intractable cases, as we will see in the next sections.
We prove Theorem 6.1 in the remaining of this section. Because the proof is technical, we sketch it here before giving the details. We first make a preliminary observation: the tractability of resilience is preserved under the mirror operation on regular languages. We then start from a word featuring a repeated letter, and pick it to optimize certain criteria (maximizing the gap between the repeated letters, and then maximizing the total length of the word): we call this a maximal-gap word, and the existence of such words is the only place where we use the fact that $L$ is finite.
@@
We first make a straightforward observation on the mirror of languages. The mirror of a word $\alpha=a_{1}\cdots a_{n}\in\Sigma^{*}$ is $\alpha^{\mathrm{R}}=a_{n}\cdots a_{1}\in\Sigma^{*}$ (in particular $\epsilon^{\mathrm{R}}=\epsilon$). The mirror of a language $L\subseteq\Sigma^{*}$ is $L^{\mathrm{R}}\coloneq\{\alpha^{\mathrm{R}}\mid\alpha\in L\}$. Observe that $(L^{\mathrm{R}})^{\mathrm{R}}=L$. It is immediate by symmetry that a language $L$ is infix-free iff $L^{\mathrm{R}}$ is infix-free. Further, it is also immediate that a language and its mirror have the same complexity for resilience:
Proposition 6.3.
@@
The point of this claim will be to eliminate symmetric cases, up to replacing the language by its mirror. We only use the claim about set semantics in the sequel (given that we will show hardness of resilience in set semantics, from which hardness for bag semantics follows): the claim for bag semantics in Proposition 6.3 is given just for completeness.
Proof of Proposition 6.3.
@@
Let $L$ be an infix-free finite language that contains a word with a repeated letter, and let us show that $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard, to establish Theorem 6.1. We will pick a word with a repeated letter for the rest of the proof, but we will choose such a word carefully. Here is the precise notion we will need:
@@
It is clear that any finite language that contains a word with a repeated letter must have maximal-gap words. Hence, in the rest of the proof, we fix $\alpha=\beta a\gamma a\delta$ to be an arbitrary such word. We will do case disjunctions on this word in the rest of the proof.
@@
Claim 6.5.
Let $L^{\prime}$ be an infix-free language containing a maximal-gap word $\beta a\gamma a\delta$ such that $\beta\neq\epsilon$ and $\delta\neq\epsilon$. Then $L^{\prime}$ is four-legged.
@@
Consider the two decompositions of the word $\alpha$ as $\alpha=(\beta a\gamma)a\delta$ and $\alpha=\beta a(\gamma a\delta)$. Note that $\beta a\gamma$, $\delta$, $\beta$, and $\gamma a\delta$ are all non-empty, so they can be used as legs. So consider the word $\alpha^{\prime}=(\beta a\gamma)a(\gamma a\delta)$. The word $\alpha^{\prime}$ contains two $a$’s separated by $\gamma a\gamma$, which is longer than $\gamma$. Hence, $\alpha^{\prime}$ is not in $L^{\prime}$, as it would contradict the fact that $\alpha$ is maximal-gap (the first criterion of the definition of maximal-gap suffices). So we have a witness of the fact that $L^{\prime}$ is four-legged. ∎
If $\beta$ and $\delta$ are both non-empty, then applying 6.5 and Theorem 5.3 with $L^{\prime}\coloneq L$ concludes. Hence, in the rest of the argument, we assume that one of $\beta$ and $\delta$ is empty. Further, without loss of generality, up to replacing $L$ by $L^{\mathrm{R}}$ using Proposition 6.3, we assume that $\beta=\epsilon$. Hence, in the rest of the proof, we can assume that $\beta=\epsilon$: we have a maximal-gap word $\alpha=a\gamma a\delta\in L$ and this word is maximal-gap.
@@
Lemma 6.6.
Let $L^{\prime}$ be an infix-free language containing a word $a\gamma a\delta$ where no infix of $\gamma a\gamma$ is in $L^{\prime}$. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is NP-hard.
Figure 7. Completed gadget for the proof of Lemma 6.6 in the case when $\delta=\epsilon$ (left), and condensed hypergraph of matches (right).
Figure 8. Completed gadget for Lemma 6.6 in the case where $\delta\neq\epsilon$ (left), and condensed hypergraph of matches (right).
@@
In the case where $\delta=\epsilon$, we use the completed gadget pictured on Figure 7 (left), where the edge labeled $\gamma$ stands for a (possibly empty) path. Note that, no matter what $\gamma$ stands for, there will be no facts having $t_{u}$ and $t_{v}$ as their head (except the two facts of the gadget completion, pictured in orange), so indeed it is a pre-gadget. The condensed hypergraph of matches is shown on Figure 7 (right), and it is an odd path, so indeed the instance pictured on Figure 7 (left) is (the completion of) a gadget and we can conclude by Proposition 4.11. Note how we use the assumption that the matches on the completed gadget are the ones indicated: directed walks including two $a$’s must include $a\gamma a$ as an infix and so they are dominated, and directed walks including only one $a$ cannot be a match of the query.
In the case where $\delta\neq\epsilon$, we use the completed gadget pictured on Figure 8 (left), again with the edge labeled $\gamma$ standing for a possibly empty path. Again, no matter what $\gamma$ stands for, there will be no facts having $t_{u}$ and $t_{v}$ as head (except for the facts added in the completion), so again we have a pre-gadget. The condensed hypergraph of matches is shown on Figure 8 (right). Now, to show that it is actually like this, first observe that the facts of the $\delta$-walks are always dominated by the preceding $a$: there is no match of the query within $\delta$ (as it is a strict infix of $a\gamma a\delta$), so all matches using $\delta$ must use the preceding $a$, and we can remove the facts of $\delta$-walks in the condensation via the node domination rule. We are then essentially back to the preceding case. More precisely, again, no directed walk containing only one copy of the $a$-edges of the gadget can be a word of the language, because these paths are either infixes of $\gamma a\delta$ (hence strict infixes of $a\gamma a\delta\in L$) or they are infixes of $\gamma a\gamma$ and so they are not in the language by assumption. Further, for any two consecutive $a$’s in the gadget, there is a $a\gamma a\delta$ match containing them. Further, after observing that $\delta$ facts are eliminated by node domination, we note that all query matches that are infixes of a word of the form $a(\gamma a)^{*}$ must contain two consecutive $a$’s, so they are subsumed by the matches corresponding to $a\gamma a\delta$, which after node domination contain just two consecutive $a$’s and the $\gamma$ between them. (The facts corresponding to $\gamma$ are then eliminated by node domination.)
Hence, we indeed have a condensed hypergraph of matches corresponding to Figure 8 (right), and we can conclude by Proposition 4.11. ∎
@@
Claim 6.7.
Let $L^{\prime}$ be an infix-free language containing the word $\alpha=a\gamma a\delta$ and containing an infix of $\gamma a\gamma$. Then $L^{\prime}$ contains a word of the form $\alpha^{\prime}=\gamma_{1}a\gamma_{2}$ where $\gamma_{1}$ is a non-empty suffix of $\gamma$ and $\gamma_{2}$ is a non-empty prefix of $\gamma$.
@@
Consider the infix of $\gamma a\gamma$ which exists by hypothesis. This infix must include the middle $a$, as otherwise it is an infix of $\gamma$, which is a strict infix of $\alpha\in L^{\prime}$, and this contradicts the assumption that $L^{\prime}$ is infix-free.